OpenStudy (anonymous):

At the beginning of an experiment, a scientist has 176 grams of radioactive goo. After 165 minutes, her sample has decayed to 5.5 grams. What is the half-life of the goo in minutes? Find a formula for G(t) , the amount of goo remaining at time t.G(t)=? How many grams of goo will remain after 50 minutes?

OpenStudy (anonymous):

ick goo

OpenStudy (anonymous):

you want to solve \[(\frac{5.5}{176})^{\frac{t}{160}}=\frac{1}{2}\] for t

OpenStudy (anonymous):

sorry should be \[(\frac{5.5}{176})^{\frac{t}{165}}=\frac{1}{2}\]

OpenStudy (anonymous):

or maybe \[0.03125^{\frac{t}{165}}=.5\] to work this with decimals. we get \[\frac{t}{165}=\frac{\ln(.5)}{\ln(.03125)}\] and so \[t=\frac{165\ln(.5)}{\ln(.03125)}\]

OpenStudy (anonymous):

whatever this is, it is your half life

OpenStudy (anonymous):

by some miracle this is a whole number, it is 33

OpenStudy (anonymous):

and so your formula is (if you don't like the one i wrote above) \[G(t)=176\times (\frac{1}{2})^{\frac{t}{33}}\] and now you can replace t by whatever number you like